OpenStudy (anonymous):

saifoo please help...complete the square -2((x^2)+3x-4)...im not sure how to factor after you do the (b/2)^2

OpenStudy (saifoo.khan):

Sorry to say i suck @ completing the square. Sat & others rules at that.

OpenStudy (anonymous):

that is for sure!! lol

OpenStudy (anonymous):

can you help me satellite73? =]

OpenStudy (anonymous):

is this \[-2(x^2+3x-4)\]?

OpenStudy (saifoo.khan):

i wanna learn that too. :D

OpenStudy (anonymous):

yes actually it started off as a transformation and then simplified to this and then it was said to find answer by completing the square

OpenStudy (saifoo.khan):

(if that gets into my head) :[

OpenStudy (anonymous):

ok why don't we start with \[-2(x^2+3x)+8\] instead.

OpenStudy (anonymous):

then half of 3 is 3/2, and 3/2 squared is 9/4 so we write \[-2(x+3x+\frac{9}{4})+8+2\times \frac{9}{4}\]

OpenStudy (anonymous):

x^2 + 3x - 4 is not a perfect square, but x^2 + 3x + (3/2)^2 is a perfect square 3/2^2 = 9/4, which is 2.25 In order to add 2.25, you have to also subtract it so that you don't change the value: -2((x^2)+3x + 2.25 - 2.25 -4) Then simplify: -2((x+3/2)^2 - 2.25-4) -2((x+3/2)^2 - 6.25)

OpenStudy (anonymous):

and the reason the last part is \[+2\times \frac{9}{4}\] is because you are actually subtracting that in the first term by the distributive law

OpenStudy (anonymous):

the original equation was f(x)=-2x^2-6x+8

OpenStudy (anonymous):

then you have \[-2(x+\frac{3}{2})^2+8+\frac{9}{2}\] or \[-2(x+\frac{3}{2})^2+\frac{25}{2}\]

OpenStudy (anonymous):

ok well then i think we have the right answer

OpenStudy (anonymous):

-2(x^2 + 3x + 9/4 - 4 - 9/4) = 0 x^2 + 3x + 9/4 = 4 + 9/4 (x + 3/2)^2 = 4 + 9/4 4(x+3/2)^2 = 16 + 9 = 25 2(x+3/2) = 5 or 2(x+3/2) = -5 2x + 3 = 5 or 2x + 3 = -5 x = 1 or x = -4

OpenStudy (anonymous):

i will write it again step by step \[f(x)=-2x^2-6x+8\] \[f(x)=-2(x^2+3x)+8\] \[f(x)=-2(x^2+3x+\frac{9}{4})+8+2\times \frac{9}{4}\] \[f(x)=-2(x+\frac{3}{2})^2+\frac{25}{2}\] and that is it

OpenStudy (anonymous):

ok ummm once again how did you get the 2x(9/4) thing...like why? and how did you factor out the (3/2)

OpenStudy (anonymous):

you did ask to complete the square, not to set it equal to zero and solve correct?

OpenStudy (anonymous):

yea i needed the complete the square

OpenStudy (anonymous):

ok first where did the 3/2 come from. the middle term is \[3x\] so i need to take half of 3, and that is \[\frac{3}{2}\]

OpenStudy (anonymous):

so once i "complete the square " my answer will look like \[(x+\text{ something })^2\] and that "something" has to be half of the coefficient of the middle term. if you multiply out you will see why

OpenStudy (anonymous):

\[x^2+3x+9/4\]is a perfect square created by completing the square, which simplifies to \[(x+\frac{3}{2})^2\]satellite got the 2x(9/4) thingy by taking (-9/4) outside of the expression in the parentheses

OpenStudy (anonymous):

now when replace \[x^2+3x\] by \[(x+\frac{3}{2})^2\] they are not the same thing

OpenStudy (anonymous):

because \[(x+\frac{3}{2})^2=x^2+3x+\frac{9}{4}\] so i have actually added \[\frac{9}{4}\]

OpenStudy (anonymous):

but in this case we had \[-2(x^2+3x)\] so when i replace that by \[-2(x+\frac{3}{2})^2\] i did not actually add \[\frac{9}{4}\] but in fact i subtracted \[2\times \frac{9}{4}\] because of that pesky -2 out front

OpenStudy (anonymous):

so in order to keep the function the same, i had to add it on back at the end. that is why i wrote \[+2\frac{9}{4}\] at the end. there is an easier way to do this however

OpenStudy (anonymous):

what is the easier way?

OpenStudy (anonymous):

nvm i got it!!! =] thnx! =]